previous 5 previous 5 of 30 NH4ci O HClo/Kolo Submit My Answers Give Up Correct Part B For the best system, calculate the ratio of the masses of the buffer components required to make the buffer Express your answer using two significant figures. m (NH3) m (NHC) Submit My Answers Give up
Answer #1

For best buffer mol of NH3 and NH4Cl should be same

Let mol of NH3 and NH4Cl be 1 mol

We know,
Mass = number of mol * molar mass

Since mol is 1 mol,
m(NH3)/m(NH4Cl) = molar mass of NH3 / molar mass of NH4Cl


Molar mass of NH3,
MM = 1*MM(N) + 3*MM(H)
= 1*14.01 + 3*1.008
= 17.034 g/mol


Molar mass of NH4Cl,
MM = 1*MM(N) + 4*MM(H) + 1*MM(Cl)
= 1*14.01 + 4*1.008 + 1*35.45
= 53.492 g/mol


So,
m(NH3)/m(NH4Cl) = molar mass of NH3 / molar mass of NH4Cl
= 17.034 / 53.492
= 0.318

Answer: 0.32